OpenStudy (anonymous):

DG's Plumbing and Heating charges $50 plus $60 per hour for emergency service. Bill remembers being billed just over $300 for an emergency call. How long to the nearest hour was the plumber at Bill's house? ____ hours

OpenStudy (zzr0ck3r):

50+60x=300

OpenStudy (zzr0ck3r):

understand how I got that?

OpenStudy (anonymous):

you added 60 and 30

OpenStudy (zzr0ck3r):

you are going to get charged 50 no matter what, then 60 bucks for each hour 1 hour 50+60(1) 2 hours 50+60(2) . . . x hours 50+60(x) and he was billed 300 so 50+60x=300 subtract 50 60x=250 divide by 60 x = 250/60

OpenStudy (zzr0ck3r):

\[x \approx 4.16\]

OpenStudy (zzr0ck3r):

and to the nearest hour is 4.0

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (anonymous):

Aw ok it does thanks for clarifying the answer